OpenStudy (anonymous):

2(b-4)= 4(2b+1) solve and check

OpenStudy (anonymous):

2b-8=8b+4 ===> b=-5/3

OpenStudy (anonymous):

i need the steps

OpenStudy (anonymous):

2(b-4)= 4(2b+1) 2b-8=8b+4 6b=-12 b=-2 now u can check by subing in b=-2 and if both sides of the equation = each other then it is correct

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

ie: -12=-12